A manufacturer of major appliances provides the following information about the operations of the refrigeration division:

Fixed costs per period are $26 880; variable costs per unit are $360; selling price per unit is $640; and capacity is 150 units.

  1. Compute

    1. the contribution margin;

    2. the contribution rate.

  2. Compute the break-even point

    1. in units;

    2. as a percent of capacity;

    3. in sales dollars.

  3. Determine the break-even point in sales dollars if fixed costs are increased to $32 200.

  4. Determine the break-even point as a percent of capacity if fixed costs are reduced to $23 808, while variable costs are increased to 60% of sales.
Answer #1

Answer:

1

the contribution margin

selling price per unit

640

Less: Variable costs per unit

360

Contribution margin per unit

280

the contribution rate.

contribution rate.

= Contribution margin per unit / sales per unit

=280/640

=43.75%

____________________________________________________

Compute the break-even point

in units;

the break-even point in units

=Fixed cost / Contribution margin per unit

=26,880 / 280

=96 units

as a percent of capacity;

=Break even point in units / total capacity

=96 /150

=64%

in sales dollars.

=Fixed cost / Contribution margin rate

=26,880 / 43.75%

=$61,440

____________________________________________________________

Determine the break-even point in sales dollars if fixed costs are increased to $32 200.

break-even point in sales dollars if fixed costs are increased to $32 200

=Fixed cost / Contribution margin rate

=32,200 / 43.75%

=$61,440

=$73,600

________________________________________________________________

Determine the break-even point as a percent of capacity if fixed costs are reduced to $23 808, while variable costs are increased to 60% of sales.

selling price per unit

640

Less: Variable costs per unit(640*60%)

384

Contribution margin per unit

256

the break-even point in units

=Fixed cost / Contribution margin per unit

=23808 / 256

=93 units

as a percent of capacity;

=Break even point in units / total capacity

=93 /150

=62%